Find parameter p so that y = 3x is a tangent line to the curve
y = x 2 + p

Grade:12

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear nitin,

  • The slope of the tangent is equal to 3. At the point of tangency y ' = 3.

    y ' = 2 x = 3

  • Solve the equation 2 x = 3 to find the x coordinate of the point of tangency.

    x = 3 / 2

  • The point of tangency is on the line y = 3 x, therefore the y coordinate of the point of tangency is equal to

    y = 3 (3/2) = 9 / 2

  • The point of tangency (3 / 2 , 9 / 2) is on the curve
    y = x 2 + p and is therefore used to find p as follows


    9 / 2 = (3 / 2) 2 + p

  • Solve the above for p to obtain

    p = 9 / 4

We are all IITians and here to help you in your IIT JEE preparation.

All the best.

 If you like this answer please approve it....

win exciting gifts by answering the questions on Discussion Forum

 

Sagar Singh

B.Tech IIT Delhi